a) In a liquid with density 1300 kg/m3, longitudinal waves with frequency 400 Hz are found to have wavelength 8.00 m. Calculate the bulk modulus of the liquid. (b) A metal bar with a length of 1.50 m has a density of 6500 kg/m3. A longitudinal sound takes 3.90 x 10-4s to travel from one end of the bar to the other. Determine the Young
Answer #1


A) Bulk Modulus B = v^2*rho

rho is the density of liquid = 1300 kg/m^3

v = lamda*f = 8*400= 3200 m/s

Bulk modulus B = (3200)^2*1300 = 1.33*10^10 Pa

B) speed of the wave v = d/t = (1.5)/(3.9*10^-4) = 3846.15 m/s

Young's modulus Y = v^2*rho = (3846.15)^2*6500 = 9.61*10^10 Pa